# Image segmentation of treated and untreated tumor spheroids by Fully Convolutional Networks

### v2

@@ Abstract
Multicellular tumor spheroids (MCTS) are advanced cell culture systems for assessing the impact of combinatorial radio(chemo)therapy. They exhibit therapeutically relevant in-vivo-like characteristics from three-dimensional (3D) cell-cell and cell-matrix interactions to radial pathophysiological gradients related to proliferative activity and nutrient/oxygen supply, altering cellular radioresponse. The therapeutic outcome of MCTS is most frequently monitored over time via brightfield microscopy imaging. Classical yet non-curative analytical readouts relate to spheroid size and volume growth kinetics. State-of-the-art assays also quantify long-term curative endpoints based on collected image time series from large treated spheroid populations per irradiation dose and treatment arm. Here, spheroid control probabilities are documented analogous to in-vivo tumor control probabilities based on Kaplan-Meier curves. These analyses require spheroid segmentation to extract relevant structural information from the images, e.g., diameter, area, volume and circularity. This laborious process entails manual inspection and analysis of up to 100.000 images per treatment arm. While several image analysis algorithms are available for spheroid segmentation, they all focus on compact MCTS with clearly distinguishable outer rim throughout growth. However, treated MCTS may partly be detached and destroyed and are usually obscured by dead cell debris. Thus, this work aims to develop an advanced automatic segmentation that also covers the previously neglected cases of spheroids obscured by debris.
Methods
We successfully train two Fully Convolutional Networks, U-Net and HRNet, and optimize their hyperparameters to develop an automatic segmentation for both untreated and treated MCTS. We systematically validate the automatic segmentation on larger, independent data sets of spheroids derived from two human head-and-neck cancer cell lines.
@@ Results
We find an excellent overlap between manual and automatic segmentation for most images, quantified by Jaccard indices at around 90%. For images with smaller overlap of the segmentations, we demonstrate that this error is comparable to the variations across segmentations from different biological experts (inter-observer variability), suggesting that these images represent biologically unclear or ambiguous cases.
@@ I Introduction
Figure 1: Representative examples of images for automatic segmentation with the optimized U-Net (blue) compared to the manual segmentation (green) for 3D tumor spheroids after treatment. The overlap with the manual segmentation is excellent for standard size and larger spheroids obscured by cell debris (top row) and sufficient for small, heavily obscured spheroids (bottom row). Displayed are (a,d) the original images, which are also the input for the U-Net, (b,d) magnified image details around the spheroids as indicated by the white box in (a,d) for visibility, and (c,e) corresponding contours from the segmentations. The metrics for evaluation of the automatic segmentation are: the top row - JCD $=7.6\%$, RDD $=3.7\%$ and RCD $=1.6\%$; bottom row - JCD$=22.7\%$, RDD$=7.5\%$ and RCD$=14.7\%$, see text for details.
@@ III.1 Evaluation metrics
RDDRCD$=\frac{|c_{P}-c_{T}|}{c_{T}}$ (2)
where $c_{P}$ is either mean diameter or circularity based on the automatic segmentation and $c_{T}$ is the corresponding characteristic calculated based on the manual segmentation. In the rare case that several spheroids are detected in an image, the largest one is chosen for computation of the JCD. In the case of JCD$=1$, which corresponds to an invalid spheroid, RDD and RCD are also set to 1. The combination of the JCD, ASF, ISF, RDD, and RCD quantify the accuracy of the segmentation with the FCN, where always a value closer to zero means higher agreement between the segmentations.
@@ IV.1 Hyperparameter optimization
For each hyperparameter set and pre-processing method a model is trained and the performance of these trained models is compared for the validation data set based on the proposed evaluation metrics from Section III.1. For JCD, RDD, and RCD not only the average value but also the standard deviation over the whole validation data set is reported to estimate reliability of the segmentation. For the ISF and the ASF the standard error of the mean of the corresponding binomial distribution is reported. All results of the optimization are listed in the SI (SI Figs. A.5-A.14) and briefly discussed in the following.
@@ IV.2 U-Net vs HRNet
The final U-Net model has a size of 158 MB and the final HRNet model a size of 251 MB. For both models, it takes about 1.8 seconds to segment one image on the CPU (Intel(R) Core(TM) i7-4770). When the segmentation is performed serially on the GPU (NVIDIA GeForce RTX 3080) the U-Net needs only 0.03 seconds per image and the HRNet 0.08 seconds. As computation time for both models is comparable, the U-Net is chosen for the automatic segmentation due to its slightly higher accuracy in out setting. Note that in a recently published deep-learning approach, the comparison of HRNet and U-Net implied that HRNet achieved the highest accuracy Lacetal2021 .
@@ IV.3 Further independent validation
Only a few images with larger spheroids exhibit JCD bigger than $0.2$. Most of these are ambiguous cases, with two spheroids attached to each other, which are inconsistently recognized as either one or two spheroids even by the human, see examples with JCD$=0.28-0.34$ and $0.42$ in App. Fig. A.4. While the training data set did not contain such double-spheroid cases, the optimized U-Net often segments them correctly, see examples with JCD$=0.08,0.12$ in SI Figs. A.2 and A.3, respectively. Apart from this particular case, most images with larger deviations between the segmentations refers to smaller spheroids with surrounding cell debris. Note that smaller spheroids without cell debris, i.e., untreated or shortly after treatment, do not exhibit such deviations but are accurately segmented. To thoroughly examine these larger deviations in more detail, three biological experts independently segmented 101 randomly selected images with spheroid diameter $d_{T}\leq 400\ \mu$m. For these 101 images the JCDs between the humans and the U-Net and between different humans are compared in Fig. 5. We find that on average the discrepancies between humans and U-Net are comparable to the variations across segmentations from different humans. This implies that these images represent biologically difficult, unclear, or ambiguous cases. It suggests that the larger JCD observed for some smaller spheroids rather reflects this uncertainty and not a low performance of the U-Net. Accordingly, the evaluation in Fig. 4 refers to a worse-case scenario, as for each spheroid, the biological expert intentionally segmented all images over time, while in practice, segmentation is only required and performed for a fraction of these images, in particular for clearly distinguishable and larger spheroids.
@@ V Discussion
Availability and requirements
Segmentation tool is published as open-source repository https://igit.informatik.htw-dresden.de/aagef650/spheroidsegdedeb (additionally download the model from https://wwwpub.zih.tu-dresden.de/~s6079592/SEGMODEL.onnx and copy it into the model/ folder) or latest version for download https://wwwpub.zih.tu-dresden.de/~s6079592/. Requirements and usage are documented in the provided README and some example images for testing are included in the corresponding folder. Training, validation, and test data used for this work will be available after peer-reviewed publication as well as the code to train and evaluate the model (see above repo training_scripts/). All networks optimized for different hyperparameters are available in the .pth format at https://wwwpub.zih.tu-dresden.de/~s6079592/models/.
Acknowledgement
@@
Funding
The authors acknowledge that this research has been co-financed by the EU, the European Social Fund (ESF), and by tax funds on the basis of the budget passed by the Saxon state parliament (project SAB-Nr. 100382145) and the Bundesministerium für Bildung und Forschung (BMBF 16dkwn001a/b). The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
Declaration of Competing Interest
The authors declare that they have no competing interests.
@@ Appendix A Supplemental figures
Table A.3: Further validation of the trained U-Net on wide variety of published test data sets from previous deep-learning models. While the trained U-Net performs well on on roughly half of the data sets or 38% of the images (with average IoU above 0.8), sometimes surpassing the original model corresponding to the data set, two types of images turn out problematic: (i) images with ambiguous ground truth (20% of images) for which the U-Net may actually segment reasonably, see main text for detailed discussion, and (ii) images on which the spheroids appears semi-transparent (42% of images), with individual cells being visible throughout the spheroid, potentially due to its small size or the chosen microscopy method. However, classical segmentation techniques work sufficiently well for both types of images (i) and (ii), making the use of deep-learning approaches in these cases unnecessary. To demonstrate this, the last two columns report the performance of the classical approach from the original publication and simple Otsu thresholding (sometimes after some image erosion for (ii)) performed by us. Note that the notation for data sets from Refs. IvaParWalAleAshGelGar2014 ; Lacetal2021 stands for brightfield/fluorescence microscopy (B/F), Nikon Eclipse/Leica DMi8/Olympus microscope (N,L,O), 2x/5x/10x magnification (2/5/10), and suspension/collagen culture (S/C). Bold values highlight the optimum for each dataset.

### v3

@@ Abstract
Multicellular tumor spheroids (MCTS) are advanced cell culture systems for assessing the impact of combinatorial radio(chemo)therapy as they exhibit therapeutically relevant in-vivo-like characteristics from 3D cell-cell and cell-matrix interactions to radial pathophysiological gradients. State-of-the-art assays quantify long-term curative endpoints based on collected brightfield image time series from large treated spheroid populations per irradiation dose and treatment arm. This analyses require laborious spheroid segmentation of up to 100.000 images per treatment arm to extract relevant structural information from the images, e.g., diameter, area, volume and circularity. While several image analysis algorithms are available for spheroid segmentation, they all focus on compact MCTS with clearly distinguishable outer rim throughout growth. However, they often fail for the common case of treated MCTS, which may partly be detached and destroyed and are usually obscured by dead cell debris.
@@ Results
To address these issues, we successfully train two Fully Convolutional Networks, UNet and HRNet, and optimize their hyperparameters to develop an automatic segmentation for both untreated and treated MCTS. We extensively test the automatic segmentation on larger, independent data sets and observe high accuracy for most images with Jaccard indices around 90%. For cases with lower accuracy, we demonstrate that the deviation is comparable to the inter-observer variability. We also test against previously published datasets and spheroid segmentations.
@@ I Introduction
Figure 1: Representative examples of images for automatic segmentation with the optimized U-Net (blue) compared to the manual segmentation (green) for 3D tumor spheroids after treatment. The overlap with the manual segmentation is excellent for standard size and larger spheroids obscured by cell debris (top row) and sufficient for small, heavily obscured spheroids (bottom row). Displayed are (a,d) the original images, which are also the input for the U-Net, (b,d) magnified image details around the spheroids as indicated by the white box in (a,d) for visibility, and (c,e) corresponding contours from the segmentations. The metrics for evaluation of the automatic segmentation are: the top row - JCD $=7.6\%$, RDD $=3.7\%$ and RCD $=1.6\%$; bottom row - ${\leavevmode\text{JCD}}=22.7\%$, RDD$=7.5\%$ and RCD$=14.7\%$, see text for details.
@@ III.1 Evaluation metrics
RDD,RCD$=\frac{|c_{P}-c_{T}|}{c_{T}}$ (2)
where $c_{P}$ is either mean diameter or circularity based on the automatic segmentation and $c_{T}$ is the corresponding characteristic calculated based on the manual segmentation. In the rare case that several spheroids are detected in an image, the largest one is chosen for computation of the JCD. This strategy is in accordance with experimental standard practice for single spheroid-based assays with curative analytical endpoints, as the largest regrowing spheroid is sufficient to detect relapse after treatment. Note that if there were several spheroids present and of interest, also all of them could be segmented. In the case of JCD$=1$, which corresponds to an invalid spheroid, RDD and RCD are also set to 1. The combination of the JCD, ASF, ISF, RDD, and RCD quantify the accuracy of the segmentation with the FCN, where always a value closer to zero means higher agreement between the segmentations.
@@ IV.1 Hyperparameter optimization
For each hyperparameter set and pre-processing method a model is trained and the performance of these trained models is compared for the validation data set based on the proposed evaluation metrics from Section III.1. For JCD, RDD, and RCD not only the average value but also the standard deviation over the whole validation data set is reported to estimate reliability of the segmentation. For the ISF and the ASF the standard error of the mean of the corresponding binomial distribution is reported. All results of the optimization are listed in the supplemental material (Figs. A.5-A.14) and briefly discussed in the following.
@@ IV.2 U-Net vs HRNet
The final U-Net model has a size of 158 MB and the final HRNet model a size of 251 MB. For both models, it takes about 1.8 seconds to segment one image on the CPU (Intel(R) Core(TM) i7-4770). When the segmentation is performed serially on the GPU (NVIDIA GeForce RTX 3080) the U-Net needs only 0.03 seconds per image and the HRNet 0.08 seconds. As computation time for both models is comparable, the U-Net is chosen for the automatic segmentation due to its slightly higher accuracy in our setting. Note that in a recently published deep-learning approach, the comparison of HRNet and U-Net implied that HRNet achieved the highest accuracy Lacetal2021 .
@@ IV.3 Further independent validation
Only a few images with larger spheroids exhibit JCD bigger than $0.2$. Most of these are ambiguous cases, with two spheroids attached to each other, which are inconsistently recognized as either one or two spheroids even by the human, see examples with JCD$=0.28-0.34$ and $0.42$ in App. Fig. A.4. While the training data set did not contain such double-spheroid cases, the optimized U-Net often segments them correctly, see examples with JCD$=0.08,0.12$ in Figs. A.2 and A.3, respectively. Apart from this particular case, most images with larger deviations between the segmentations refers to smaller spheroids with surrounding cell debris. Note that smaller spheroids without cell debris, i.e., untreated or shortly after treatment, do not exhibit such deviations but are accurately segmented. To thoroughly examine these larger deviations in more detail, three biological experts independently segmented 101 randomly selected images with spheroid diameter $d_{T}\leq 400\ \mu$m. For these 101 images the JCDs between the humans and the U-Net and between different humans are compared in Fig. 5. We find that on average the discrepancies between humans and U-Net are comparable to the variations across segmentations from different humans. This implies that these images represent biologically difficult, unclear, or ambiguous cases. It suggests that the larger JCD observed for some smaller spheroids rather reflects this uncertainty and not a low performance of the U-Net. Accordingly, the evaluation in Fig. 4 refers to a worse-case scenario, as for each spheroid, the biological expert intentionally segmented all images over time, while in practice, segmentation is only required and performed for a fraction of these images, in particular for clearly distinguishable and larger spheroids.
@@ V Discussion
VI Availability of source code and requirements (optional, if code is present)
  • Project name: Spheroidsegdedeb (Spheroid segmentation despite debris)
  • Operating system(s): Platform independent
  • Programming language: Python ¿3.8.10
  • Other requirements: Additionally download the model from https://wwwpub.zih.tu-dresden.de/~s6079592/SEGMODEL.onnx. Requirements and usage are documented in the provided README and some example images for testing are included in the corresponding folder.
  • License: GNU GPL
  • BioTool ID: spheroidsegdedeb
  • SciCrunch ID: SpheroidSegDeDeb, RRID:SCR_026409
  • SpheroidSegDeDeb is also archived in Software Heritage snapshot .
VII Data availability
Training, validation, and test data used for this work are available bioimage including data on clean spheroids and the extended dataset as well as the code to train and evaluate the model (see above repo training_scripts/). All additional supporting data are available in the GigaScience repository, GigaDB gigadb , i.e. all networks optimized for different hyperparameters in the .pth format. DOME-ML annotations are available in DOME registry dome .
VIII Declarations
VIII.1 List of abbreviations
Ambiguous Spheroid Fraction (ASF), fully convolutional network (FCN), Invalid Spheroid Fraction (ISF), Jaccard Distance (JCD), Multicellular tumor spheroids (MCTS), Relative Circularity Deviation (RCD), Relative Diameter Deviation (RDD), Supplemental material (SM)
VIII.2 Competing Interests
The authors declare that they have no competing interests.
VIII.3 Funding
The authors acknowledge that this research has been co-financed by the EU, the European Social Fund (ESF), and by tax funds on the basis of the budget passed by the Saxon state parliament (project SAB-Nr. 100382145) and the Bundesministerium für Bildung und Forschung (BMBF 16dkwn001a/b). The publication of this article is funded by the Open Access Publication Fund of Hochschule für Technik und Wirtschaft Dresden – University of Applied Sciences. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
VIII.4 Author’s Contributions
Conceptualization, A.V.-B., L.A.K.-S., and S.L.; methodology, A.V.-B., S.L., S.M., L.A.K.-S., and M.S..; software, M.S., K.L., and S.L.; validation, W.C., K.L., S.M., and S.L.; formal analysis, M.S., A.V.-B., and S.L.; investigation, M.S., K.L., and S.L.; resources, L.A.K.-S., and A.V.-B.; data curation, M.S., W.C., K.L., and S.L.; writing – original draft preparation, M.S., A.V.-B. and S.L.; writing – review and editing, A.V.-B., W.C., S.L., S.M., and L.A.K.-S.; visualization, M.S., W.C., and S.L.; supervision, L.A.K.-S., A.V.-B. and S.L.; project administration, L.A.K.-S. and A.V.-B.; funding acquisition, L.A.K.-S. and A.V.-B. All authors have read and agreed to the published version of the manuscript.
IX Acknowledgements
@@ Appendix A Supplemental figures
Table A.3: Further validation of the trained U-Net on wide variety of published test data sets from previous deep-learning models. While the trained U-Net performs well on on roughly half of the data sets or 38% of the images (with average IoU above 0.8), sometimes surpassing the original model corresponding to the data set, two types of images turn out problematic: (i) images with ambiguous ground truth (20% of images) for which the U-Net may actually segment reasonably, see main text for detailed discussion, and (ii) images on which the spheroids appears semi-transparent (42% of images), with individual cells being visible throughout the spheroid, potentially due to its small size or the chosen microscopy method. However, classical segmentation techniques work sufficiently well for both types of images (i) and (ii), making the use of deep-learning approaches in these cases unnecessary. For demonstration, the last two columns report the performance of the classical approach from the original publication and simple Otsu thresholding (sometimes after some image erosion for (ii)) performed by us for images of types (i) and (ii). Note that the notation for data sets from Refs. IvaParWalAleAshGelGar2014 ; Lacetal2021 stands for brightfield/fluorescence microscopy (B/F), Nikon Eclipse/Leica DMi8/Olympus microscope (N,L,O), 2x/5x/10x magnification (2/5/10), and suspension/collagen culture (S/C). Bold values highlight the optimum for each dataset.